OpenStudy (goformit100):

In three coloured boxes - Red, Green and Blue, 108 balls are placed. There are twice as many balls in the green and red boxes combined as there are in the blue box and twice as many in the blue box as there are in the red box. How many balls are there in the green box ? A. 18 B. 36 C. 45 D. None of these

OpenStudy (goformit100):

@lilminiwin

OpenStudy (unicornpoopcookies):

Answer: Option D Explanation: Let R, G and B represent the number of balls in red, green and blue boxes respectively. Then, . R + G + B = 108 ...(i), G + R = 2B ...(ii) B = 2R ...(iii) From (ii) and (iii), we have G + R = 2x 2R = 4R or G = 3R. Putting G = 3R and B = 2R in (i), we get: R + 3R + 2R = 108 6R = 108 R = 18. Therefore Number of balls in green box = G = 3R = (3 x 18) = 54.

OpenStudy (goformit100):

Thank this helped me very much.

OpenStudy (unicornpoopcookies):

No problem. :)